We are working with rref matrices. what are the possible solutions to these matrices?

7. Describe all solutions to: [ 2 -2 [ 4 0 1 1 101 -9 21 | T = [2010] 14 i 21] [3] ſo 3 2 0 0] 3 3 2 2 ·ī=
Answer #1

what are the matries? possible salution to theise 12 0107 -2 1-9 14 1 21 [2 0 107 1:2 2 2 2 whe ran دم is [ 2o1o7 1-2 1-4 l.nRank of coeffient marrin - Rank of Augmented matrin = 2 & The eqution are consistent or salveble. But Rank of coeffient makriT 2010 - 21-gã=o [4121] Augmented matrin is 2 o 10:17 7- 21 ago 14 1 21 31 R2 = R₂ + R, [2 o 10: 17 o i Rz? = R₂_2R l R, =R_RFrom equin (3) n +1 21 in ₂t mal >) 1,21-m . The saletion of makin is n = 1-10m ₂ a 1-m ng -m wherer m EZ2 (m is an integn 10Feran I & From ear o ol Wit P+ q = 4. Miz 4-p-q These motrint They the salitions 1 i n i of mahrin is ,= 4-p-g n 2 2 q na a p

Like please